# Emergent Matryoshka doll-like point gap in a non-Hermitian quasiperiodic lattice

### v1

@@ V Conclusion
In summary, we study localization and topological phase transitions in a GSM quasiperiodic lattice. The results reveal that multiple MEs and point gaps with high winding number can emerge under the condition of summation terms’ number $N>1$. The system’s PT symmetry is damaged synchronously with the phase transition. Specifically, for the case of finite summation number $N$, the circle number of topological point gaps can at most be $N$. As the eigenstates continuously transform into localized states, the different point gaps gradually merge, and the maximum winding number gradually increases from $\omega=0$ to $\omega=N$. For the limit case of $N=\infty$, however, Avila’s global theory can give the exact ME’s expression. Analytical and numerical results consistently show that the corresponding MEs (topological point gaps) will merge together, resulting in a winding number $\omega=1$. Furthermore, the main results confirm that the properties of MEs and point gap structures are consistent with the behavior of topological boundaries and PT symmetry breaking.
VI Acknowledgements
@@
We thank R.-J. Chen, X.-D. Hu and W.-L. Li for their insightful suggestions. This work was supported by the National Key Research and Development Program of China (Grant No.2022YFA1405300), and the Guangdong Basic and Applied Basic Research Foundation (Grants No.2021A1515012350).
@@
Figure 5: The detialed evolution of point gap and the corresponding winding number with repsect to $\lambda$ under the condition of $N=2$. The qusiperiodic strength are marked. The other parameters $L=610$, $\alpha=377/610$ and $q=0.2$.
@@
Figure 6: The detialed evolution of point gap and the corresponding winding number with repsect to $\lambda$ under the condition of $N=4$. The qusiperiodic strength are marked. The other parameters $L=610$, $\alpha=377/610$ and $q=0.05$.
Appendix A The evolution of point gap with respect to $\lambda$
@@
To exhibit the Matryoshka doll-like point gaps and the process of finding the base point $E_{b}$, we provide two examples, i.e., the case of $N=2$ and $N=4$.
@@
First, let’s discuss the case of $N=2$. It is not difficult to find from Fig. 5 that with the increase of quasiperiodic strength, some states with the largest eigenenergy in the system undergo localization phase transitions first. Then, as $\lambda$ continues to increase, some eigenstates with the lowest energy begin to enter the localized state. Then, the number of localized states will continue to increase with an increasing $\lambda$, and eventually all extended states will tranform into localized states. Then, the two points will merge together to form a two-layer Matryoshka doll-like structure. After that, by further increasing $\lambda$, the size of the ring become even larger, and the two-layer circile structure will always exist.
@@
Now we turn to the case of $N=4$. With the increase of quasiperiodic strength $\lambda$, the pace of phase transition is simular to the former case, i.e., the states in the two ends of eigenenergy localized first, and the states in the middle will follow suit. Besides, unlike the $N=2$ case, there are more point gaps with higher-order winding number. Specifically, there are point gaps with winding number $\omega=1$ (e.g., subfigures of $\lambda=0.4$), $\omega=2$ (e.g., subfigures of $\lambda=1.0$), point gaps of $\omega=3$ (e.g., subfigures of $\lambda=1.4$), $\omega=4$ (e.g., subfigures of $\lambda=1.6$) [see Fig. 6].
Appendix B Discussions on phase diagram of different $N$
@@
Figure 7: $\Gamma$ phase diagram on $\lambda-E$ plane with different $N$. Parameters $L=610$, $\alpha=377/610$ and $q=0.2$.
@@
The relationship between the ME and the parameter $N$ can be summarized as follows. Under the condition of $N=1$, the system exhibits a traditional topological point gap, in which there is no ME and only a standard critical point, namely $\lambda_{c}=J$. With the increase of parameter $N$, multiple MEs gradually appear in the system, and the number of MEs is $2(N-1)$, which is consistent with the previous prediction. Finally, as $N$ increases, the exponential decay term at large $N$ makes the MEs all merge together, while only one ME can be observed on the phase diagram. One can obtain the same conclusion from the point of view of topological point gap analysis, i.e., although the increase of point gaps’s circle promotes a higher winding number with an increasing $N$, the maximum winding number decreases to $1$ due to the merging of point gaps’ circles.

### v2

@@ VI Conclusion
In summary, we study localization properties and topological point gaps in a GSM quasiperiodic lattice. The results reveal that multiple MEs and point gaps with high winding number can emerge under the condition of summation terms’ number $N>1$. The system’s PT symmetry is damaged synchronously with the topological phase transition. Specifically, for the case of finite summation number $N$, the circle number of topological point gaps can at most be $N$. As the eigenstates continuously transform into localized states, the different point gaps gradually merge, and the maximum winding number gradually increases from $\omega=0$ to $\omega=N$. For the limit case of $N=\infty$, however, Avila’s global theory can give the exact ME’s expression. Analytical and numerical results consistently show that the corresponding MEs (topological point gaps) will merge together, resulting in a winding number $\omega=1$. Furthermore, the main results confirm that the properties of MEs and point gap structures are consistent with the behavior of topological boundaries and PT symmetry breaking. In principle, the corresponding phenomena can be realized by artificial quantum systems. As an example, we propose an experimental scheme based on the Rydberg atomic array (see Appendix F).
VII Acknowledgements
@@
We thank R.-J. Chen, X.-D. Hu and W.-L. Li for their insightful suggestions. This work was supported by the National Key Research and Development Program of China (Grant No.2022YFA1405300), Open Fund of Key Laboratory of Atomic and Subatomic Structure and Quantum Control (Ministry of Education) and the Guangdong Basic and Applied Basic Research Foundation (Grants No.2021A1515012350).
Appendix A Details for GSM lattice model
To make a better understanding of the model, we provide deduction details. The Hamiltonian reads
$\begin{split}H=&\sum_{j=1}^{L-1}(Jc_{j}^{\dagger}c_{j+1}+\mathrm{H.c.})\\ &+\sum_{j=1}^{L}\sum_{n=1}^{N}{\lambda}e^{-(n-1)q}e^{in(2{\pi}{\alpha}j+{\theta})}c_{j}^{\dagger}c_{j},\end{split}$ (18)
where
$V_{j}=\lambda e^{i(2\pi\alpha j+\theta)}$ (19)
for $N=1$;
$V_{j}={\lambda}e^{i(2{\pi}{\alpha}j+{\theta})}+\lambda e^{-q}e^{2i(2{\pi}{\alpha}{j}+\theta)}$ (20)
for $N=2$;
$V_{j}={\lambda}e^{i(2{\pi}{\alpha}j+{\theta})}+{\lambda}e^{-q}e^{2i(2{\pi}{\alpha}j+{\theta})}+{\lambda}e^{-2q}e^{3i(2{\pi}{\alpha}j+{\theta})}$ (21)
for $N=3$.
The terms number of GSM potential increases with an increasing $N$. The potential forms a geometric sequence. Interestingly, as $N\rightarrow\infty$, the terms number can reduce to be one by using the geometric sequence formula. Then, one can obtain
$V_{j}=\frac{\lambda{e}^{i(2{\pi}{\alpha}j+{\theta})}-\lambda{e}^{-{N}q}{e}^{({N}+1)i(2{\pi}{\alpha}{j}+{\theta})}}{1-{e}^{-q}e^{i(2{\pi}{\alpha}j+{\theta})}}.$ (22)
Note that, when $N\rightarrow\infty$, the term $\lambda{e}^{-{N}q}{e}^{({N}+1)i(2{\pi}{\alpha}{j}+{\theta})}$ tends to 00. then, one can obtain
Thus, the terms number is one and the corresponding maximum winding number reduces to $\omega=1$. By the way, Eq. (23) is a non-Hermitian version of the Ganeshan-Pixley-Das Sarma potential SGaneshan2015 .
@@ Appendix A Details for GSM lattice model
Figure 6: The detailed evolution of point gap and the corresponding winding number with respect to $\lambda$ under the condition of $N=2$. The quasiperiodic strengths are marked. The other parameters $L=610$, $\alpha=377/610$ and $q=0.2$.
Appendix B The evolution of point gap with respect to $\lambda$
@@
Figure 7: The detailed evolution of point gap and the corresponding winding number with respect to $\lambda$ under the condition of $N=4$. The quasiperiodic strength are marked. The other parameters $L=610$, $\alpha=377/610$ and $q=0.05$.
@@
To exhibit the multi-loop nested point gaps and the process of finding the base point $E_{b}$, we provide two examples, i.e., the case of $N=2$ and $N=4$.
@@
First, let’s discuss the case of $N=2$. It is not difficult to find from Fig. 6 that with the increase of quasiperiodic strength, some states with the largest eigenenergy in the system undergo localization phase transitions first. Then, as $\lambda$ continues to increase, some eigenstates with the lowest energy begin to enter the localized state. Then, the number of localized states will continue to increase with an increasing $\lambda$, and eventually all extended states will transform into localized states. Then, the two points will merge together to form a two-layer multi-loop nested structure. After that, by further increasing $\lambda$, the size of the ring become even larger, and the two-layer circle structure will always exist.
@@
Now we turn to the case of $N=4$. With the increase of quasiperiodic strength $\lambda$, the pace of phase transition is similar to the former case, i.e., the states in the two ends of eigenenergy localized first, and the states in the middle will follow suit. Besides, unlike the $N=2$ case, there are more point gaps with high winding number. Specifically, there are point gaps with winding number $\omega=1$ (e.g., subfigure of $\lambda=0.4$), $\omega=2$ (e.g., subfigure of $\lambda=1.0$), point gaps of $\omega=3$ (e.g., subfigure of $\lambda=1.4$), $\omega=4$ (e.g., subfigure of $\lambda=1.6$) [see Fig. 7].
Appendix C Discussions on phase diagram of different $N$
@@
Figure 8: $\Gamma$ phase diagram on $\lambda-E$ plane with different $N$. Parameters $\lambda=1.9$, $L=610$, $\alpha=377/610$ and $q=0.2$.
@@
The relationship between the ME and the parameter $N$ can be summarized as follows. Under the condition of $N=1$, the system exhibits a traditional topological point gap, in which there is no ME and only a standard critical point, namely $\lambda_{c}=J$. With the increase of parameter $N$, multiple MEs gradually appear in the system, and the number of MEs is $2(N-1)$, which is consistent with the previous prediction. Finally, as $N$ increases, the exponential decay term at large $N$ makes the MEs all merge together, while only one ME can be observed on the phase diagram. One can obtain the same conclusion from the point of view of topological point gap analysis, i.e., although the increase of point gaps’ circle promotes a higher winding number with an increasing $N$, the maximum winding number decreases to $1$ due to the merging of point gaps’ circles.
Appendix D Scaling effect
According to the definition Eq. (3), the fractal dimension is related to the system size. To ensure the correction, this section provides the result of finite size scaling in Fig. 9. As shown in the figure, the system size affects the points density of the loop structure in the complex energy plane, but does not qualitatively affect the localization and topological characteristics of the system. In concrete terms, the point distribution is sparse due to the small size [see the case of $L=89$]. The energy spectrum is smoother and more complete with increasing calculation points. The calculation accuracy increases accordingly. Once the size exceeds $L=610$, neither the smoothness nor the completeness of the image undergoes further changes. Thus, $L=610$ is the optimal choice considering computing power constraints.
Figure 9: The finite size scaling of point gap with $N=2$ (a), $N=3$ (b) and $N=\infty$ (c). The system size $L$ are marked. The corresponding $\alpha=\frac{55}{89},~{}\frac{89}{144},~{}\frac{144}{233},~{}\frac{377}{610}$ and $\frac{1597}{2584}$ respectively. Throughout, $q=0.2$.
Figure 10: The point gap with $N=2$ (a) and $N=3$ (b) tuned by parameter $q$. (c) The phase diagram of winding number ${\omega}$ in the $\lambda-q$ plane when $N=2$. The other parameters $\alpha=377/610$ and $L=610$.
Figure 11: Experimental scheme on GSM model in a Rydberg atomic array. (a) Schematic diagram of atomic energy levels. (b) Rydberg atoms trapped in optical tweezers to form the atomic array.
Appendix E The effect of $q$
The parameter $q$ acts on the exponent of Eq. (2) and plays an important role in modulating the quasiperiodic strength. In concrete terms, $q\to\infty$, $e^{-(n-1)q}\to 0$. On the other word, when $q$ increases, the quasiperiodic strength vanishes and the system becomes trivial regardless of the parameters adjusted. Only the metallic phase is presented, whereas the Anderson localization disappears. Therefore, the system becomes trivial for a large $q$. So, we consider $q\in(0,1)$ in our discussion. We provide the cases of $N=2,3$ as examples [see Fig. 10]. When $q$ increases to $1$, extended states emerge from the system whose states were originally all localized [see Fig. 10(a)]. Accordingly, the topological phase of the system gradually changes from $\omega=2$ to $\omega=1$. The results of $N=3$ are similar [see Fig. 10(b)].
Furthermore, we calculate a phase diagram of the winding number $\omega$ in $\lambda-q$ plane. The system undergoes a topological phase transition with an increasing $\lambda$ [see Fig. 10(c)]. More importantly, the topological phase transition point is delayed as $q$ increases. The system fails to achieve $\omega=2$ until $q$ reaches a certain threshold. This is consistent with the previous results. For the purpose of effective discussion, we set $q=0.2$.
Appendix F Experimental realization and measurement
The model we propose in this paper is an extended version of the AA model. So far, the standard AA model has been realized in a variety of artificial systems, such as optical lattice ultracold atoms, ion traps, optical waveguide arrays, superconducting lines, Rydberg atomic array, and so on. The manipulation techniques of quasiperiodic disorder in these artificial quantum simulators are very mature. Based on the existing experimental platforms of AA model, it only needs to calculate the specific GSM potential at the corresponding lattice sites by strictly following the potential function expression given in this paper, i.e., Eq. (2) in the main text.
Following this, as an example, we discuss the relevant experimental scheme on the 87Rb Rydberg array. Since the Rydberg atomic array corresponds to a spin model, we need to map the Rydberg atomic array to the model described by Eq. (1). Experimentally, the Hamiltonian required for the Rydberg experiment is given by
$H_{s}=\sum_{j_{x}}(J\sigma_{j_{x}}^{+}\sigma_{j_{x}+1}^{-}+H.c.)+\frac{1}{2}\sum_{j_{x}}V_{j_{x}}(\mathbb{I}+\sigma_{j_{x}}^{z}),$ (24)
where $\sigma^{\pm}=\frac{1}{2}(\sigma_{x}\pm i\sigma_{y})$ with Pauli matrices $\sigma_{x}$ and $\sigma_{y}$. The Hamiltonian (24) can be transformed to the GSM lattice by defining operator $c^{\dagger}_{j}=\sigma^{+}_{j_{x}}=\left|\uparrow\right>_{j_{x}}\left<\downarrow\right|_{j_{x}}$ at each site $j_{x}$, where |$=70P_{1/2}$ and |$=70S_{1/2}$ for 87Rb atoms [see Fig. 11]. The dipole-dipole interaction between Rydberg atoms is given by $J=\frac{d^{2}}{R^{3}}$, where $d$ represents the transition dipole moment between the two Rydberg levels, $R$ is the distance between two nearest neighboring sites, $70$ is the principal quantum number, and the energy unit $J=2\pi\times 7.53$MHz. Note that, since the non-nearest neighboring effect decays with distance $R^{\nu}$ ($\nu\geq 3$), one can safely ignore the corresponding terms.
Specifically, the experiment on Rydberg atomic array includes four steps. The first step is to load the atoms into the optical tweezers. The second step is to excite the atoms to the Rydberg state of the principal quantum number $70$. The third step is to introduce the GSM potential energy into the atomic array. The fourth step is to measure the eigenvalues and eigenstates of the system and then calculate the corresponding fractal dimension. A detailed introduction will be provided below.
At first, optical tweezers, which are generated by a spatial light modulator (SLM), will be used to form an atomic chain of 87Rb atoms [see Fig. 11]. Specifically, one can capture the atoms by optical tweezers, and then rearrange them by an acousto-optical deflector (AOD). Consequently, one can obtain a one-dimensional array of 87Rb atoms.
Secondly, the 87Rb atoms will be excited from the $5S_{1/2}$ state to the $70S_{1/2}$ state. One can achieve it by coupling the $5S_{1/2}$ and $70S_{1/2}$ states through 420nm and 1012nm lasers, with the $6P_{1/2}$ state acting as the intermediate state during this two-photon process.
Thirdly, by irradiating a 1012nm Rydberg laser to the corresponding atoms (with a detuning of about 100MHz to the 1012nm Rydberg laser) and considering the decoherence TMWintermantel2020 , different sites can achieve an on-site potential, which satisfies the quasiperiodic properties. This potential is given by $V_{AC}=V_{on-site}=\frac{|\Omega_{0}|^{2}}{\Delta}[\frac{1}{2}\sum_{n=1}^{N}e^{-(n-1)q}e^{in(2{\pi}{\alpha}j_{x})}]^{2}$, where $\Omega_{0}$ is the corresponding laser Rabi frequency. Specifically, this process can be carried out by introducing a second spatial light modulator (SLM). By generating a locally controllable light shift, one can achieve site-dependent detuning $\Delta_{\text{eff},n}$, as marked in Fig. 11(b). The corresponding method has recently been reported by the Harvard group TManovitz2024 .
At last, one can use a dynamical method to measure the corresponding phenomena on the Rydberg array PRoushan2017 ; SZLi2025 . In concrete terms, one can excite a Rydberg atom to a superposition state of $70S_{1/2}$ and $70P_{1/2}$ as the initial state, and then let the system evolve. The eigenvalues and the corresponding density distributions can be obtained by Fourier transform.